# Article Category

### v1

@@ 5.2 BOUNDARY PROJECTION
with safe action set $\mathcal{A}_{s}=\left<c_{\mathcal{A}_{s}},G_{\mathcal{A}_{s}}\right>$. Translating Equation 17 to canonical, quadratic form yields the optimisation variable $z=\begin{bmatrix}a_{s}\\ \gamma\end{bmatrix}\in\mathbb{R}^{d+n}$ and the problem
@@
where the subscript of the identity denotes its size, and bold scalars are matrices of appropriate size filled with the scalar. To compute the rank of the Jacobian, we start from the differentials of the KKT conditions [DifferentiateQuadraticProgram, Eq. 6]
@@
where the subscript $D$ denotes a diagonalised vector, the superscript * optimal values, $\nu\in\mathbb{R}^{d}$ the dual variables on the equality constraints, $\lambda\in\mathbb{R}^{n}$ the dual variables on the inequality constraints, and $\mathop{}\!\mathrm{d}$ a differential. To obtain the Jacobian with respect to the action, we substitute $\mathop{}\!\mathrm{d}q=\frac{\mathop{}\!\mathrm{d}q}{\mathop{}\!\mathrm{d}a}\mathop{}\!\mathrm{d}a=-2\begin{bmatrix}I_{d}\\ \textbf{0}\end{bmatrix}\mathop{}\!\mathrm{d}a$ and all other differential terms with zero leaving
@@
$\frac{\mathop{}\!\mathrm{d}\lambda}{\mathop{}\!\mathrm{d}a}-G_{\mathcal{A}_{s}}^{T}\frac{\mathop{}\!\mathrm{d}\nu}{\mathop{}\!\mathrm{d}a}$ $=\textbf{0}$ (29)
$\lambda_{D}^{*}\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}+(\gamma_{D}^{*}-I_{n})\frac{\mathop{}\!\mathrm{d}\lambda}{\mathop{}\!\mathrm{d}a}$ $=\textbf{0}$ (30)
$\frac{\mathop{}\!\mathrm{d}a_{s}}{\mathop{}\!\mathrm{d}a}-G_{\mathcal{A}_{s}}\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}$ $=\textbf{0}\,.$ (31)
We solve this system by substitution starting from Equation 31
@@
which we substitute into Equation 28
@@
We substitute this into Equation 29
@@
with which we decouple Equation 30
@@
The constraint is inactive $\lambda_{i}^{*}=0$ or active $\gamma_{i}^{*}=1$ to fulfil the conditions. We define the index set of active constraints as $\mathcal{I}_{a}=\left\{i\mid\gamma_{i}^{*}=1\right\}$ and of inactive constraints as $\mathcal{I}_{i}=\left\{i\mid\lambda_{i}^{*}=0\right\}$. We examine the rows of Equation 35
@@
$\lambda_{i}^{*}\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{i,:}$ $=0$ $i\in\mathcal{I}_{a}$ (38)
@@
We partition the Jacobian according to the index sets $\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}=\begin{bmatrix}\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{a},:}\\ \left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{i},:}\end{bmatrix}$. The assumption of strong duality implies strict complementarity $\lambda_{i}^{*}>0$. Therefore, Equation 38 specifies the submatrix $\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{a},:}=\textbf{0}$. The left-hand side of Equation 39 reduces to
@@
using Equation 38. We isolate the remaining submatrix
@@
$\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{a},:}$ $=\textbf{0}$ (42)
@@
We plug this result into Equation 32
@@
The resulting matrix is the projection matrix onto the column space of $\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}$. It is also unique since any homogeneous part of the solution lies in the null space of $\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}$, see Section 7.6 for details.

### v2

@@ 7.6 PROOF OF LEMMA 1: JACOBIAN OF BOUNDARY PROJECTION
with safe action set $\mathcal{A}_{s}=\left<c_{\mathcal{A}_{s}},G_{\mathcal{A}_{s}}\right>$. Translating Equation˜40 to canonical, quadratic form yields the optimisation variable $z=\begin{bmatrix}a_{s}\\ \gamma\end{bmatrix}\in\mathbb{R}^{d+n}$ and the problem
@@
where the subscript of the identity denotes its size, and bold scalars are matrices of appropriate size filled with the scalar. To compute the rank of the Jacobian, we start from the differentials of the KKT conditions [76, Eq. 6]
@@
where the subscript $D$ denotes a diagonalised vector, the superscript * optimal values, $\nu\in\mathbb{R}^{d}$ the dual variables on the equality constraints, $\lambda\in\mathbb{R}^{n}$ the dual variables on the inequality constraints, and $\mathop{}\!\mathrm{d}$ a differential. To obtain the Jacobian with respect to the action, we substitute $\mathop{}\!\mathrm{d}q=\frac{\mathop{}\!\mathrm{d}q}{\mathop{}\!\mathrm{d}a}\mathop{}\!\mathrm{d}a=-2\begin{bmatrix}I_{d}\\ \textbf{0}\end{bmatrix}\mathop{}\!\mathrm{d}a$ and all other differential terms with zero, leaving
@@
[…]
We solve this system by substitution starting from Equation˜54
@@
which we substitute into Equation˜51
@@
We substitute this into Equation˜52
@@
with which we decouple Equation˜53
@@
The constraint is inactive $\lambda_{i}^{*}=0$ or active $\gamma_{i}^{*}=1$ to fulfil the conditions. We define the index set of active constraints as $\mathcal{I}_{a}=\left\{i\mid\gamma_{i}^{*}=1\right\}$ and of inactive constraints as $\mathcal{I}_{i}=\left\{i\mid\lambda_{i}^{*}=0\right\}$. We examine the rows of Equation˜58
@@
$\lambda_{i}^{*}\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{i,:}$ =0=0= 0 $i\in\mathcal{I}_{a}$ (61)
@@
We partition the Jacobian according to the index sets $\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}=\begin{bmatrix}\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{a},:}\\ \left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{i},:}\end{bmatrix}$. The assumption of strong duality implies strict complementarity $\lambda_{i}^{*}>0$. Therefore, Equation˜61 specifies the submatrix $\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{a},:}=\textbf{0}$. The left-hand side of Equation˜62 reduces to
@@
using Equation˜61. We isolate the remaining submatrix
@@
[…]
@@
We plug this result into Equation˜55
@@
The resulting matrix is the projection matrix onto the column space of $\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}$. It is also unique since any homogeneous part of the solution lies in the null space of $\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}$, see Section˜7.7 for details.